### v2

@@ 5.5 How do TSLMs work with persistent series?
As described above, some of the series are transformed before entering the estimation stage of the models. Looking at the results shown so far, there appears to be some indication that TSLMs perform relatively better when the degree of persistence of the series they utilize is relatively low. To further investigate this aspect, we break down the previous results by separately looking at series with different degrees of persistence. Specifically, Figure 7 shows results for variables that enter the models in levels or log levels, and are therefore relatively persistent, while Figure 8 shows results for variables that enter the model in log-differences, yielding typically a less persistent variable.
@@
As one would expect, BVARs perform relatively better when estimated with more persistent variables (likely because keeping the variables in levels allows for common trends and co-integration) while DFMs perform relatively better when estimated with less persistent variables (likely because the estimation of the factors relies on stationarity). This is evident in the graphs, where the BVARs box-plots are consistently below 1 in Figure 7 and the DFMs box-plots are consistently below 1 in Figure 8. Turning the attention to TSLMs, they seem to perform better when applied to stationary variables (even though it should be noted that they are still outperformed by DFMs for this subset of series). When TSLMs encounter persistent variables their performance is more mixed, and is systematically worse at the 1-step ahead horizon.
@@
Figure 7: Distribution of RMSFEs (relative to an AR benckmark) for the econometric models and a selection of TSLMs, focusing on variables with transformation codes equal to 1 or 4. The evaluation sample is January 1985 to December 2019.
@@
Figure 8: Distribution of RMSFEs (relative to an AR benckmark) for the econometric models and a selection of TSLMs, focusing on variables with transformation codes equal to 5. The evaluation sample is January 1985 to December 2019.
6 Conclusions and directions for further research
In this paper, we have presented a comparative analysis evaluating the forecasting accuracy of Time Series Language Models against state of the art econometric forecasting models. The analysis has been performed with the widely used FRED-MD database. We have found that the forecasting performance of some TSLM is comparable to that of the BVARs and factor models, but the latter appear to be generally more reliable, as they tend to be more robust to structural changes and less prone to generate the occasional unreasonable forecast. Some caveats should be kept in mind. First, the TSLMs we considered were pre-trained by developers on datasets of their choice, might or might not contain information on the variables they are set to forecast, and since the training data lacks timestamps, it’s hard to reproduce the situation of a forecaster in real time. Second, the TSLMs we have considered here have all been used with a zero-shot implementation. This means that there is room to improve their performance via fine tuning. Addressing these issues is part of our ongoing research.

### v3

@@ 5.5 How do TSLMs Work with Persistent Series?
As described above, some of the series are transformed before entering the estimation stage of the models. Looking at the results shown so far, there appears to be some indication that TSLMs perform relatively better when the degree of persistence of the series they utilize is relatively low. To further investigate this aspect, we break down the previous results by separately looking at series with different degrees of persistence. Specifically, Figure 6 shows results for variables for which the one lag partial autocorrelation is below 0.9, while Figure 7 shows results for variables for which the one lag partial autocorrelation is above 0.9. The latter variables are of course highly persistent.
@@
For the econometric models, the pattern emerging from this particular cut of the results is in line with expectations. Specifically, VARs perform relatively better when estimated with more persistent variables (likely because keeping the variables in levels allows for common trends and co-integration) while factor models perform relatively better when estimated with less persistent variables (likely because the estimation of the factors relies on stationarity). This is evident in the graphs, where the factor model box-plots are consistently below 1 in Figure 6 while the BVAR box-plots are consistently below 1 in Figure 7.
@@
Figure 6: Distribution of RMSFEs (relative to an AR benckmark) for the econometric models and a selection of TSLMs, focusing on variables with medium and low persistence. The evaluation sample is January 1985 to December 2019.
@@
Figure 7: Distribution of RMSFEs (relative to an AR benckmark) for the econometric models and a selection of TSLMs, focusing on variables with high persistence. The evaluation sample is January 1985 to December 2019.
Turning the attention to TSLMs, for series with low to moderate persistence, they seem to perform relatively better, even though it should be noted that they are still outperformed by factor models for this subset of series. The overall picture changes starkly when considering highly persistent variables. In these cases, one can observe a definite increase in the variation of the performance of TSLM.
To give an example of the extent of variation in performance, Figure 8 shows the time series of out-of-sample forecasts produced by the factor model, the BVAR (with symmetric priors) and TimesFM model for a few selected variables. While for the 1-step ahead horizons the forecasts are quite similar (left side panels), large differences emerge at the 12-step ahead horizon (right side panels).
Figure 8: Time series of out-of-sample forecasts produced by the factor model, the BVAR with symmetric priors and the timesFM model, for selected variables. Left panels show 1-step ahead forecasts, and right panels show 12-step ahead forecasts. The sample is January 1985 to December 2019.
For Housing Starts (HOUST), the 12-step ahead forecasts produced by TimesFM are remarkably good, as this model is able to almost perfectly predict the pronounced slowdown that happened during the 2007 financial crisis and the ensuing recession. This results in an astonishing 50 percent improvement in RMSFE compared to the AR benchmark. Instead, the econometric models only pick up this slowdown with a delay, and the factor model even fails to recognize the depth of the trough. Of course, one must keep in mind that TimesFM includes this recessionary time period in its training sample, so it seems to be on some level aware that a change in time series behavior is to be expected within these dates. For weekly hours (CES0600000007) TimesFM outperforms the factor model but underperforms the BVAR. For variables such as the spread between the 10 year interest rate and the Federal Funds rate (T10YFFM) and the spread between corporate bonds and the Federal Funds rate (BAAFFM), the performance of TSLMs is subpar, featuring entire periods of "hallucination" in the second half of the 80s and in 2017-2019, resulting in RMSFE losses of the order of 15-20 percent.
Figure 9: Distribution of RMSFEs (relative to an AR benckmark) for TimesFM and Moirai-large for zero shot and fine-tuning. The evaluation sample is January 1985 to December 2019.
6 Fine-tuning
Up until now, all TSLMs results have been based on zero-shot predictions of the original pretrained models. As discussed in subsection 2.2, alternatively one could consider the option of fine-tuning the TSLMs, i.e. updating the pretrained model’s parameters on a task-specific dataset, with the goal of improving forecasting performance. Broadly speaking, to implement fine-tuning one could follow one of two approaches, (i) Continued-pretraining or (ii) Parameter-efficient fine-tuning. [^17] See Han et al. (2024) for a detailed review. Continued-pretraining involves further training the model on a domain-specific dataset starting with pretrained parameters, and allowing all model parameters to adapt to the characteristics of the given data. This can lead to overfitting when the fine-tuning data is small. Parameter-efficient methods instead freeze some of the original model’s parameters and introduce trainable low-rank matrices (e.g. low-rank adaptation (LoRA) methods as in (Hu et al., 2022)), significantly reducing the number of trainable parameters, addressing overfitting, and requiring comparatively less computational resources.
Here, we selected the two best models from our zero-shot experiments – TimesFM and Moirai-large – and fine-tuned them on FRED-MD data. Note that some of the pretrained TSLMs, including Moirai, are trained on datasets that overlap with the FRED-MD. However, the influence of any single dataset in the pretraining phase gets diluted within the vast and diverse pretraining corpus. Consequently, the models learn general time series patterns, rather than specializing in the specific characteristics of FRED-MD. To make a TSLM more dataset-specific, fine-tuning is required. The main challenge we face in this context is that FRED-MD is a relatively small dataset, which may lead to overfitting problems. [^18] Kumar et al. (2022) discusses how in these situations the model may also lose the knowledge acquired during pretraining, leading to the so-called catastrophic forgetting. To address these challenges, we employ a simple parameter-efficient fine-tuning strategy as opposed to continued-pretraining. In particular, we freeze the transformer layers of the two TSLMs, thereby reducing the number of trainable parameters. [^19] Kumar et al. (2022) refers to this as Linear Probing. Note that freezing the pretrained model’s weights (parameters) is a special case of LoRA, which additionally injects trainable rank decomposition matrices into each layer of TSLM. After freezing the transformer layers, the number of trainable parameters reduced to approximately $\leq 10$M for both TimesFM (instead of $200$M) and Moirai-large (instead of $311$M), primarily part of the final layers. [^20] In the case of Moirai, we also consider fine-tuning using the continued-pretraining strategy, as this was the recommended approach provided by the authors in the model’s repository, but in this case the model ended up severely overfitting and producing poor out-of-sample forecasts. Throughout, to keep computational costs manageable, we chose to fine-tune each model once every 12 months, and the resulting models were then used for zero-shot forecasting over the subsequent 12 months. [^21] Within each fine-tuning round, the training FRED-MD data was split into an 80:20 ratio, with 20% reserved for validation. We trained these models for 200 epochs, utilizing early stopping with a patience of 5 to mitigate overfitting. To further counter overfitting and ensure stable training on our limited dataset, we reduced the learning rate, preventing large weight updates. We run the fine-tuning experiment on a stock server equipped with a single NVIDIA A6000 GPU. In terms of wall-clock time, fine-tuning took about 10 hours for each of the two models. Hence, comparatively, fine-tuning remains significantly more resource-intensive than the BVARs and factor models we considered, which can be trained on standard CPUs in a fraction of the time. [^22] We also considered the continued-pretraining option for the same models, and in that case the total computing time was approximately 90 hours.
Figure 9 shows the comparison of zero-shot and fine-tuned performance for TimesFM and Moirai-large. We observe that fine-tuning yields results that are similar or slightly better than the zero-shot performance for both models. This result overall isn’t particularly surprising for our setup: a large model and a small dataset. While fine-tuning offers the potential for adapting models to a domain, it comes with its own set of challenges, such as hyperparameter optimization and increased computational resource (e.g. GPUs) requirements, thus needs careful consideration. Given the marginal performance gains achieved through fine-tuning in our context, the inherent "plug-and-play" simplicity of zero-shot forecasting presents itself as a potentially more practical approach.
7 Conclusions
In this paper, we have provided a comprehensive evaluation of the forecasting performance of Time Series Language Models (TSLMs) in predicting macroeconomic variables. While recent advances in machine learning have led to the development of these models as promising alternatives to traditional econometric methods, our results indicate that their performance remains mixed. Specifically, only two of the five models we evaluated, Salesforce’s Moirai and Google’s TimesFM, consistently outperform a simple autoregressive benchmark. Furthermore, these same models struggle to deliver consistently superior forecasts compared to established macroeconomic forecasting methods such as Bayesian Vector Autoregressions (BVARs) and Factor Models. Our findings therefore suggest that while TSLMs can offer valuable insights, they are not yet a clear replacement for state-of-the-art econometric models.
A key challenge in applying TSLMs to macroeconomic forecasting is the lack of control over their training data. For example, many of these models are pretrained on the FRED-MD dataset, hence already contain the macroeconomic series that are the focus of our forecasting exercise, introducing potential biases and making it difficult to conduct a clean pseudo out-of-sample forecasting exercise. Furthermore, practical challenges and computational constraints make retraining these models in real time a non-trivial task. Taken together, these issues raise important concerns about the role of TSLMs in economic forecasting, particularly in comparison to existing time series methods such as BVARs and factor models which have proven extremely flexible and successful in these same settings.
Despite these limitations, our results indicate that TSLMs exhibit some promising features, particularly in capturing nonlinearities and adapting to evolving economic conditions. Their relatively stronger performance in the post-Covid-19 period suggests that they may have advantages in handling structural breaks and rapidly changing economic environments. As research in this area continues to evolve, it may be worthwhile exploring hybrid approaches that integrate the flexibility and generality of foundational models with the theoretical rigor of econometric methods. For example, it would be interesting to investigate whether incorporating structured economic constraints into the TSLMs’ architectures could enhance their reliability. We leave this topic to further research.